OpenStudy (anonymous):

HOW DO i FIND THE ANSWER Y-5=1/4 (X-2)

OpenStudy (anonymous):

I am assuming you want it in y = mx + b form y - 5 = 1/4(x - 2) Distribute the 1/4 through the parenthesis y - 5 = 1/4x - 1/2 Now add 5 to each side y - 5 + 5 = 1/4x - 1/2 + 5 Combine like terms y = 1/4x - 1/2 + 5 y = 1/4x -1/2 + 10/2 y = 1/4x + 9/2